9x+2y=26
1.5x-2y=13

Answer by TimothyLamb(4379)

9x+2y=26
1.5x-2y=13
---
put the system of linear equations into standard form
---
9x + 2y = 26
1.5x - 2y = 13
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 3.7142857
y = -3.7142857
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php